# ETH Zurich

### v1

@@ Abstract
Annotation and classification of legal text are central components of empirical legal research. Traditionally, these tasks are often delegated to trained research assistants. Motivated by the advances in language modeling, empirical legal scholars are increasingly turning to prompting commercial models, hoping that it will alleviate the significant cost of human annotation. Despite growing use, our understanding of how to best utilize large language models for legal tasks remains limited. We conduct a comprehensive study of 260 legal text classification tasks, nearly all new to the machine learning community. Starting from GPT-4 as a baseline, we show that it has non-trivial but highly varied zero-shot accuracy, often exhibiting performance that may be insufficient for legal work. We then demonstrate that a lightly fine-tuned Llama 3 model vastly outperforms GPT-4 on almost all tasks, typically by double-digit percentage points. We find that larger models respond better to fine-tuning than smaller models. A few tens to hundreds of examples suffice to achieve high classification accuracy. Notably, we can fine-tune a single model on all 260 tasks simultaneously at a small loss in accuracy relative to having a separate model for each task. Our work points to a viable alternative to the predominant practice of prompting commercial models. For concrete legal tasks with some available labeled data, researchers are better off using a fine-tuned open-source model.
@@ 1 Introduction
The legal system generates a staggering volume of complex documents. United States federal courts alone process hundreds of thousands of cases a year, each having substantial case files. Much empirical legal research involves the systematic collection and analysis of such data in order to understand how laws function in practice and what impact they have on society. What limits researchers across the board is the cost of annotating and classifying legal documents. Legal classification tasks vary in complexity, but often require substantial expertise and effort. Employing trained research assistants stretches to a few hundred, perhaps a few thousand documents at a time, but is no match for the sheer scale of legal data.
Motivated by the rapid advances in large language models, law scholars increasingly try out commercial models, such as GPT-4, on a variety of legal tasks, hoping to boost the efficiency of legal research. For lack of an alternative, researchers access these models in a zero-shot manner [^1] Fine-tuning GPT-4 is currently unavailable to all API users and requires special access.. The underlying assumption is that prompting GPT-4 provides the best solution to the problem that is currently available. In this work, we critically examine this assumption.
@@ 1.1 Our contributions
We introduce and study a collection of 260 legal classification tasks, nearly all new to the machine learning community. The tasks we introduce are actual legal classification tasks based on the U.S. Supreme Court (Spaeth et al., 2023) and Court of Appeals (Songer, ) databases that provide rich annotations for court files. We use these annotations as labels to derive challenging multi-class machine learning tasks. In particular, our suite of tasks significantly extends and strengthens valuable recent efforts to benchmark language models for legal tasks (Guha et al., 2023).
Our primary finding is that fine-tuning a single Llama 3 8B Inst (MetaAI, 2024) model on all classification tasks achieves vastly superior performance to GPT-4222We evaluate gpt-4-0613, which is what at the time of writting gpt-4 points to in the OpenAI API. The recently released GPT-4o model is currently not available for our region via the Azure OpenAI Service. zero-shot (Figure 1). Although it is expected that fine-tuning helps, the strong superiority of fine-tuning an open source model at much smaller scale is surprising. After all, GPT-4 is orders of magnitude larger according to available information and interpretation. Based on our comprehensive evaluation, we argue that the zero-shot performance of GPT-4 is far from sufficient for actual legal work. In contrast, a fine-tuned model provides a simple and practical alternative with much higher accuracy. In more detail, our primary contributions the following:
Figure 1: The cost of zero-shot: Difference in accuracy between Lawma 8B and GPT4. Each vertical bar represents the accuracy difference on one task, sorted in ascending order.
  • To establish baselines, we perform a zero-shot evaluation of recent state-of-the-art open source models (Figure 3), including Llama 3 8B Inst, Llama 3 70B Inst, Mistral 7B Inst, Saul 7B Inst, and Mixtral 8x7B Inst, and the commercial model GPT-4. Among these, only Llama 3 70B Inst and GPT-4 achieve zero-shot performance significantly better than a constant classifier that outputs the majority class. Still, there are dozens of tasks were both models perform worse than random guessing (Figure 4). Averaged across all tasks, GPT-4 has zero-shot accuracy 62.9% compared with 58.4% for Llama 3 70B. The constant classifier that outputs the majority class for each task achieves 41.7% accuracy, a number determined by class imbalance and the number of classes in each task. GPT-4 and Llama 3 achieve similar performance on the Court of Appeals tasks (63.4% and 60.3%, respectively), whereas GPT-4 has significantly higher accuracy on the Supreme Court tasks (59.8% versus 47.1%). Few-shot prompting GPT-4 does not improve performance. Section 2.3 provides more details.
  • Fine-tuning a single Llama 3 8B Instruct model, which we call Lawma 8B, outperforms GPT-4 zero-shot an all but a few of tasks. The improvements are typically in the double digit accuracy points, specifically, 22.6 percentage points on average for Supreme Court tasks and 16.5 points on average over Appeals Court tasks. Figure 1 displays the accuracy deficit of zero-shot prompting relative to fine-tuning across all tasks. In absolute terms, Lawma 8B achieves 82.4% accuracy on Supreme Court tasks and 79.9% accuracy on Appeals Court tasks.
    We additionally fine-tune a single Llama 3 70B Instruct model, which we call Lawma 70B, and obtain further small single digit improvements in performance. Lawma 70B outperforms Lawma 8B on average by 1.7 percentage points for Supreme Court tasks and 1.6 points for Appeals Court tasks.
@@
  • We systematically evaluate the effect of finetuning on model performance at different scales. Larger models respond better to fine-tuning than smaller models. Across nine different base models, mean task accuracy increases steadily with pretraining compute as Figure 7 illustrates. However, we observe signs of diminishing returns. For Appeals Court tasks, scaling fine-tuning from Pythia 1B to Llama 3 70B yields an average improvement of only 8.5 accuracy points, despite a 3000x increase in pretraining compute. This suggests that, in the future, major improvements may not come from model scale alone.
@@
  • Fine-tuning is data efficient. A few hundred examples typically suffice to achieve high classification accuracy. This is crucial, since labeling a few hundred data points is often financially feasible for many legal scholars (Hall and Wright, 2008), whereas labeling many thousands may not. Figure 8 shows how training accuracy increases with the number of training examples.
@@
  • We can simultaneously fine-tune on all 260 tasks. There is not a large loss compared with fine-tuning on a specific task (Figure 9). This is desirable in practice, as it obviates the need to train and maintain a separate model for each task.
  • Fine-tuning generalizes to unseen tasks. We show that fine-tuning Llama 3 8B Inst only on the Court of Appeals tasks improves its average accuracy on Supreme Court tasks by 18.8 accuracy points (Figure 10).
@@
  • We contextualize our accuracy numbers with intercoder agreement rates. Our analysis reveals task heterogeneity in the relationship between model accuracy and intercoder agreement.
Our results speak to the power of specialization for legal classification tasks. They challenge the prevailing narrative about the suitability of the zero-shot paradigm. Fine-tuning provides a simple, practical, and far more accurate solution than zero-shot prompting. These insights suggest that the empirical legal community should invest in an ecosystem of fine-tuned models for relevant classification tasks. From a benchmarking perspective, our tasks provide non-trivial measurements of model performance that may be of independent interest.
1.2 Related work and background
@@
The legal community has moved relatively quickly in adopting GPT models. Several startups have begun using incorporating large language models, including GPT, into legal products (Wiggers, 2022). Lexis Nexis, a major commercial provider of law-related services, has partnered with Open AI and Anthropic to offer legal text generation (LexisNexis, 2023). Legal scholars have evaluated GPT’s performance on the bar exam (Katz et al., 2024) as well as law school exam (Choi et al., 2023). Choi and Schwarcz (2023) examined how GPT-4 can improve student performance on law school exams. Nay et al. (2024) examined how LLMs perform on answering multiple choice questions related to tax law. Gray et al. (2024) used GPT models to extract information from cases concerning the factors that predict the constitutionality of police stops. Choi (2023) used GPT-4 to extract information concerning interpretative techniques from U.S. Supreme Court decisions. Livermore et al. (2023) tested the performance of GPT models for categorizing cases by issue areas and in recommending citations based on case similarity. Savelka and Ashley (2023) evaluate the zero-shot performance of GPT-4 on a variety of semantic legal annotation tasks. Engel and Mcadams (2024) ask GPT for the ordinary meaning of statutory terms. In the area of corporate law, Frankenreiter and Talley (2024) use GPT-4 to extract information about the contents of corporate charters.
Benchmarks and models for legal tasks.
@@
LegalBench (Guha et al., 2023) is a recent multi-task benchmark for natural language understanding in legal domains. As of writing, LegalBench consists of 162 tasks gathered from 40 contributors. LegalBench draws on numerous earlier benchmarking efforts in different legal domains, specifically, inference on contracts (Koreeda and Manning, 2021; Hendrycks et al., 2021), merger agreement understanding (Wang et al., 2023), identifying the legal holding of a case (Zheng et al., 2021), statutory reasoning (Holzenberger and Van Durme, 2021), privacy compliance and policy (Wilson et al., 2016; Zimmeck et al., 2019; Ravichander et al., 2019), and identifying unfair clauses in terms of service (Lippi et al., 2019). We extend and strengthen these valuable efforts to benchmark large language models in legal settings. The 260 tasks we study are complementary to those in LegalBench. We focus on core legal classification tasks based on the U.S. Supreme Court Database (Spaeth et al., 2023) and the U.S. Courts of Appeals database (Songer, ), relevant to the field empirical legal studies. These databases provide rich information about judicial decisions, panel compositions, judge attributions, and case characteristics, that we use to derive challenging classification tasks. Section 2 continues with this discussion in more detail. We did not evaluate our model on LegalBench, since our model is specialized to the Supreme Court and Appeals Court data.
Legal-BERT (Chalkidis et al., 2020) is a BERT-like transformer model that was pretrained on a few hundred thousand legal documents, including more than 100,000 documents from EU legislation, and more than 150,000 documents from the Case Law Access Project portal. SaulLM (Colombo et al., 2024) is a more recent model trained on a legal corpus of 30 billion tokens. SaulLM is based on the Mistral 7B model (Jiang et al., 2023) that incorporates additional pretraining and instruction fine-tuning specific to legal text. Our work is rather orthogonal, evaluating the strength of specialized models rather than building new general-purpose models.
@@
Hall and Wright (2008) provide an overview of the use of human annotators in empirical legal studies. Student coders have been deployed to extract a wide variety of features from legal data. Although student researchers are much less expensive than private attorneys, the costs can quickly become prohibitive. Depending on the size of the document and the complexity of the task, research assistants can label roughly dozens of examples per hour. Projects involving the labeling of hundreds of documents are financially feasible for many legal scholars, but projects involving many thousands of documents are largely impractical. In an example of a larger annotation effort, Frankenreiter et al. (2021) employed human coders to annotate several thousands of corporate charters. Using ChatGPT for a similar task, Frankenreiter and Talley (2024) estimated that employing human coders would have been approximately ten times more costly than their approach.
Data annotation and labeling also play a major role in machine learning benchmarks and applications, see, e.g., Aroyo and Welty (2015); Gray and Suri (2019); Hardt and Recht (2022) for background. Dorner and Hardt (2024) give an extended discussion about label quality and annotator disagreement in the context of machine learning benchmarks.
@@ 1.3 Limitations
Fine-tuning increases accuracy to about $80\%$ in our evaluation suite compared with around $60\%$ for zero-shot prompting. While we are rather certain that $60\%$ accuracy is insufficient for consequential legal work, we emphasize that $80\%$ is still far from perfect. In addition, the variance in accuracy across tasks, while lower than for zero-shot prompting, remains high. Although our work meets the ethical and technical recommendations by Kapoor et al. (2024) for “developers of legal AI”, we maintain caution about the use of large language models for consequential legal tasks. To which extent these models are suitable for use in specific applications requires additional substantive investigation. We add that the legal documents we consider are exclusively from either the U.S. Supreme Court or appellate courts in the United States. We cannot speak to how these results may change for tasks in other legal domains within United States or legal systems in other countries.
2 Legal classification tasks
@@
In this work, we focus on legal classification tasks. Legal classification tasks range in complexity, from extremely simple tasks that require little specialized knowledge, to highly sophisticated tasks that involve substantial expertise and judgment. Simple tasks would include identifying the parties to a case or the general issue area—for example, whether a case dealt with family law or commercial contracts. More sophisticated tasks involve specific legal knowledge, familiarity with legal principles or discourse, and the ability to engage in nuanced analogical or conceptual reasoning. For example, labeling the ideological valence of a decision requires the annotator to understand how specific legal issues map onto contemporary political debates, while labeling the standard of review applied by an appellate court requires detailed knowledge of these standards as well as the ability to parse procedural history. Many legal doctrines are quite complicated, involving multipart tests, nuanced exceptions, and balancing inquiries. Extracting features concerning such doctrines can lead to disagreement even among experienced annotators with considerable legal expertise.
@@
More efficient ways to solve legal classification tasks would be tremendously useful in practice. A well functioning system to automatically extract relevant features from legal texts could, in particular, facilitate empirical legal study across a wide range of domains. This research could include not only social scientific study of the causes or consequences of judicial decisions, but also more traditional research modalities based on doctrinal interpretation (Livermore and Rockmore, 2019). There is an almost unlimited variety of features that legal scholars could study, ranging from the factors cited by judges when deciding the outcomes of property law disputes to the relationship between the party affiliation of judges and their use of different interpretative styles. With the digitization of legal texts at the U.S. state level and outside the United States, low-cost and flexible featurization can also boost efforts to show the geographic diffusion of legal concepts or approaches.
@@
To summarize, our reasons to study legal classification tasks are both technical and substantive. From a technical machine learning perspective, these tasks provide highly non-trivial classification problems where even the best models leave much room for improvement. From a substantive legal perspective, efficient solutions to such classification problems have rich and important applications in legal research, as we discussed.

### v2

@@ Abstract
Annotation and classification of legal text are central components of empirical legal research. Traditionally, these tasks are often delegated to trained research assistants. Motivated by the advances in language modeling, empirical legal scholars are increasingly turning to prompting commercial models, hoping that it will alleviate the significant cost of human annotation. Despite growing use, our understanding of how to best utilize large language models for legal annotation remains limited. To bridge this gap, we introduce CaselawQA, a benchmark comprising 260 legal annotation tasks, nearly all new to the machine learning community. We demonstrate that commercial models, such as GPT-4.5 and Claude 3.7 Sonnet, achieve non-trivial yet highly variable accuracy, generally falling short of the performance required for legal work. We then demonstrate that small, lightly fine-tuned models outperform commercial models. A few hundred to a thousand labeled examples are usually enough to achieve higher accuracy. Our work points to a viable alternative to the predominant practice of prompting commercial models. For concrete legal annotation tasks with some available labeled data, researchers are likely better off using a fine-tuned open-source model. Code, datasets, and fine-tuned models are available at https://github.com/socialfoundations/lawma.
@@ 1 Introduction
The legal system generates a staggering volume of complex documents. United States federal courts alone process hundreds of thousands of cases a year, each having substantial case files. Much empirical legal research involves the systematic collection and analysis of such data to understand how laws function in practice and what impact they have on society. What limits researchers across the board is the cost of annotating and classifying legal documents. Legal classification tasks vary in complexity, but often require substantial expertise and effort. Employing trained research assistants stretches to a few hundred, perhaps a few thousand documents at a time, but is no match for the sheer scale of legal data.
There has long been an interest by empirical legal scholars in NLP tools for feature extraction (i.e., annotation) in lieu of human annotators (Livermore and Rockmore, 2019). Starting from sentiment analysis and topic models, to now large language models. The costs and error of existing methods is the single most important bottleneck in the empirical legal studies pipeline. Yet, the use of large language models to annotate legal text remains a critically understudied area.
Nonetheless, motivated by the rapid advances in large language models, law scholars increasingly try out commercial models, such as GPT-4, on a variety of annotation tasks, hoping to boost the efficiency of legal research (Choi, 2023; Gray et al., 2024; Frankenreiter and Talley, 2024). The underlying assumption is that large commercial models provide the best solution to the problem that is currently available. In this work, we critically examine this assumption.
@@ 1.1 Our contributions
We introduce and study a collection of 260 legal classification tasks, nearly all new to the machine learning community. The tasks we introduce are actual legal classification tasks based on the U.S. Supreme Court (Spaeth et al., 2023) and Court of Appeals (Songer, ) databases. These legal databases offer rich annotations for court cases, which we utilize as labels to create challenging multi-class classification tasks. We aggregate these tasks into an easy-to-use benchmark, which we call CaselawQA (Section 2).
Figure 1: The power of specialization: Performance of various language models on the CaselawQA benchmark for legal annotation. The Lawma models, specialized for legal annotation, outperform all other models.
Our primary finding is that, for legal annotation tasks, small fine-tuned models substantially outperform large commercial models (Figure 1). Specifically, we fine-tune on a series of language models, ranging from 135M to 70B parameters, on the CaselawQA legal annotation tasks. We collectively refer to these models as the Lawma family of models. Notably, our smallest fine-tuned model, Lawma 135M, surpasses all commercial models, achieving 82% accuracy on CaselawQA, compared to 78% for Claude 3.7 Sonnet. Our largest model, Lawma 70B, achieves 88% accuracy, outperforming every commercial model by at least 10 percentage points.
Although it is expected that fine-tuning improves performance, the superiority of fine-tuning an open-weights model at a much smaller scale is surprising. After all, the commercial models evaluated are up to four orders of magnitude larger than the smallest Lawma model. Our results demonstrate that, for concrete legal annotation tasks with some available labeled data, researchers are likely better off using small specialized models rather than large general-purpose LLMs. We conduct various large-scale fine-tuning experiments that further demonstrate the benefits and practicality of specializing models for legal annotation:
@@
  • Larger models respond better to fine-tuning. Accuracy of the Lawma models increases steadily with model size (Figure 9). However, we observe signs of diminishing returns. This suggests that, in the future, major improvements in fine-tuned performance may not come from model scale alone.
@@
  • Fine-tuning is data efficient. A few hundred to a thousand examples typically suffice to achieve higher accuracy than commercial models (Section 4.2, Figure 10). This is crucial, since labeling a few hundred examples is often financially feasible for legal scholars, whereas labeling many thousands may not.
  • Fine-tuning on the precise annotation tasks of interest is crucial. While fine-tuning Llama 3.1 8B Instruct exclusively on Court of Appeals tasks increases its accuracy on Supreme Court tasks by 14 points, it still falls short of Lawma 8B by 25 points (Section 4.3, Figure 11). This highlights how much potential performance we might be missing by not specializing precisely for the target tasks of interest.
@@
  • We can simultaneously fine-tune on all 260 annotation tasks. There is not a large loss compared with fine-tuning on a specific annotation task (Section 4.4, Figure 13). This is desirable in practice, as it obviates the need to train and maintain a separate model for each annotation task.
@@
  • We contextualize our accuracy numbers with intercoder agreement rates. Our analysis reveals task heterogeneity in the relationship between model accuracy and intercoder agreement (Section 4.5).
Our results speak to the power of specialization for legal annotation. Our insights suggest that the empirical legal community should invest in an ecosystem of fine-tuned models for relevant annotation tasks. Such an ecosystem could radically expand the capacity of legal scholars to engage in quantitative work.
From a benchmarking perspective, the tasks presented in this work are of independent interest. They are challenging multi-class classification problems that require some amount of legal expertise. The best models achieve non-trivial, but modest performance. And even fine-tuned models don’t reach intercoder agreement rates. These legal classification tasks are diverse, non-trivial evaluation tasks for future model advances.
Finally, our work challenges the prevailing narrative about the suitability of “generalist” models. In commercial APIs, users are generally limited to prompting generalist models, as fine-tuning is costly for the model provider. But as we show, generalist models are neither sufficiently good nor best possible for many practical tasks. Specializing models to concrete tasks of interests, even with relatively small base models and few labeled examples, can provide a simple, practical, and more accurate solution.
1.2 Related work
@@
The legal community has moved relatively quickly in adopting GPT models. Several startups have begun using incorporating large language models, including GPT, into legal products (Wiggers, 2022). Lexis Nexis, a major commercial provider of law-related services, has partnered with Open AI and Anthropic to offer legal text generation (LexisNexis, 2023). Legal scholars have evaluated GPT’s performance on the bar exam (Katz et al., 2024) as well as law school exam (Choi et al., 2023). Choi and Schwarcz (2023) examined how GPT-4 can improve student performance on law school exams. Nay et al. (2024) examined how LLMs perform on answering multiple choice questions related to tax law. Gray et al. (2024) used GPT models to extract information from cases concerning the factors that predict the constitutionality of police stops. Choi (2023) used GPT-4 to extract information concerning interpretative techniques from U.S. Supreme Court decisions. Livermore et al. (2023) tested the performance of GPT models for categorizing cases by issue areas and in recommending citations based on case similarity. Savelka and Ashley (2023) evaluate the zero-shot performance of GPT-4 on a variety of semantic legal annotation tasks. Engel and Mcadams (2024) ask GPT for the ordinary meaning of statutory terms. In the area of corporate law, Frankenreiter and Talley (2024) use GPT-4 to extract information about the contents of corporate charters. Before large language models, Hausladen et al. (2020) used shallow classifiers over n-grams to predict the ideological direction of Court of Appeals cases.
Benchmarks for legal tasks.
@@
LegalBench (Guha et al., 2023) is a recent multi-task benchmark for natural language understanding in legal domains. As of writing, LegalBench consists of 162 tasks gathered from 40 contributors. LegalBench draws on numerous earlier benchmarking efforts in different legal domains, specifically, inference on contracts (Koreeda and Manning, 2021; Hendrycks et al., 2021), merger agreement understanding (Wang et al., 2023), identifying the legal holding of a case (Zheng et al., 2021), statutory reasoning (Holzenberger and Van Durme, 2021), privacy compliance and policy (Wilson et al., 2016; Zimmeck et al., 2019; Ravichander et al., 2019), and identifying unfair clauses in terms of service (Lippi et al., 2019). Bhambhoria et al. (2024) evaluate the performance of general-purpose models on legal question-answering tasks and advocate for the development of open-source models specifically tailored to the legal domain. We extend and strengthen these valuable efforts to benchmark large language models in legal settings. We focus on core legal classification tasks based on the U.S. Supreme Court Database (Spaeth et al., 2023) and the U.S. Courts of Appeals database (Songer, ), which are relevant to the field of empirical legal studies. Our evaluation suite measures the performance of models in annotating court opinions, focusing on features that are of practical interest to the field of empirical legal studies. The tasks we study are complementary to those in LegalBench. We do not evaluate our model on LegalBench, since our models are specialized to the Supreme Court and Appeals Court data.
Large language models for the legal domain.
General-purpose language models are likely to be trained on a substantial amount of legal data because much of this data is publicly available on the internet. For example, the FreeLaw dataset includes a large collection of court opinions (Gao et al., 2021). Legal-BERT (Chalkidis et al., 2020) is a BERT-like transformer model that was pretrained on a few hundred thousand legal documents. The more recent SaulLM models (Colombo et al., 2024b, a) adapt the open-weights Mistral (Jiang et al., 2023, 2024) models to the legal domain both by continual pretraining and instruction-tuning on legal text. In contrast to Lawma, we consider SaulLM to be a general-purpose model for the legal domain, not tailored to any specific legal task. Our approach differs significantly; we focus on developing models specialized for annotation tasks of practical interest to empirical legal studies. We demonstrate that this specialization is highly effective, with our specialized Lawma models significantly outperforming all other evaluated LLMs.
@@
Hall and Wright (2008) provide an overview of the use of human annotators in empirical legal studies. Student coders have been deployed to extract a wide variety of features from legal data. Although student researchers are much less expensive than private attorneys, the costs can quickly become prohibitive. Depending on the size of the document and the complexity of the task, research assistants can label roughly dozens of examples per hour. Projects involving the labeling of hundreds of documents are financially feasible for many legal scholars, but projects involving many thousands of documents are largely impractical. In an example of a larger annotation effort, Frankenreiter et al. (2021) employed human coders to annotate several thousands of corporate charters. Using ChatGPT for a similar task, Frankenreiter and Talley (2024) estimated that employing human coders would have been approximately ten times more costly than their approach. Data annotation and labeling also play a major role in machine learning benchmarks and applications, see, e.g., Aroyo and Welty (2015); Gray and Suri (2019); Hardt and Recht (2022) for background. Dorner and Hardt (2024) give an extended discussion about label quality and annotator disagreement in the context of machine learning benchmarks.
@@ 1.3 Limitations
While our fine-tuned models substantially outperform commercial models, we emphasize that our fine-tuned models are still far from perfect, and the variance in accuracy across tasks remains high. Although our work meets the ethical and technical recommendations by Kapoor et al. (2024) for “developers of legal AI”, we maintain caution about the use of large language models for consequential legal tasks. To which extent these models are suitable for use in specific applications requires additional substantive investigation. We add that the legal documents we consider are exclusively from either the U.S. Supreme Court or appellate courts in the United States. We cannot speak to how these results may change for tasks in other legal domains within the United States or legal systems in other countries.
2 CaselawQA: a benchmark for legal annotation
@@
Legal annotation tasks range in complexity, from extremely simple tasks that require little specialized knowledge, to highly sophisticated tasks that involve substantial expertise and judgment. Simple tasks would include identifying the parties to a case or the general issue area—for example, whether a case dealt with family law or commercial contracts. More sophisticated tasks involve specific legal knowledge, familiarity with legal principles or discourse, and the ability to engage in nuanced analogical or conceptual reasoning. For example, labeling the ideological valence of a decision requires the annotator to understand how specific legal issues map onto contemporary political debates, while labeling the standard of review applied by an appellate court requires detailed knowledge of these standards as well as the ability to parse procedural history. Many legal doctrines are quite complicated, involving multipart tests, nuanced exceptions, and balancing inquiries. Extracting features concerning such doctrines can lead to disagreement even among experienced annotators with considerable legal expertise.
@@
More efficient ways to solve legal classification tasks would be tremendously useful in practice. A well functioning system to automatically extract relevant features from legal texts could, in particular, facilitate empirical legal study across a wide range of domains. This research could include not only social scientific study of the causes or consequences of judicial decisions, but also more traditional research modalities based on doctrinal interpretation (Livermore and Rockmore, 2019). There is an almost unlimited variety of features that legal scholars could study, ranging from the factors cited by judges when deciding the outcomes of property law disputes to the relationship between the party affiliation of judges and their use of different interpretative styles. With the digitization of legal texts at the U.S. state level and outside the U.S., low-cost and flexible featurization can also boost efforts to show the geographic diffusion of legal concepts.
@@
To summarize, our reasons to study legal classification tasks are both technical and substantive. From a technical machine learning perspective, these tasks provide highly non-trivial classification problems where even the best models leave much room for improvement. From a substantive legal perspective, efficient solutions to such classification problems have rich and important applications in legal research.